### v2

# Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis

Abstract
CP violation in baryon decays has recently garnered significant attention. Traditionally, CP violation in charmed baryon decays is predicted to be at the order of $O(10^{-4})$. Utilizing recent measurements from BESIII, Belle and Belle II experiments, we conduct a comprehensive global analysis to estimate the contribution of topological diagrams in these processes. Our study reveals that topological diagrams alone cannot fully account for the experimental data. By incorporating new effects that are not included in traditional topological diagrams and exploring a novel way to determine the possible new weak phase, our analysis reveals the potential for significant CP violation in $\Lambda_{c}^{+}\to p\eta/n\pi^{+}$ and $\Xi_{c}^{+}\to\Xi^{0}K^{+}$. This suggests a promising opportunity to observe CP violation for the first time in charmed baryon decays.
preprint: CTPU-PTC-24-18
Introduction- CP violation (CPV) is one of the most important issues in particle physics Lee:1957qs ; Kobayashi:1973fv ; Deshpande:1994ii ; LHCb:2019hro ; Lenz:2020awd ; Schacht:2021jaz ; Bediaga:2022sxw ; Wang:2022tcm ; Wang:2022fih ; Maccolini:2023luv ; Shen:2023eln ; Song:2024jjn ; He:2024pxh ; Schacht:2024wgc ; Wang:2024ztg , necessary to explain the baryon-antibaryon asymmetry of the universe. However, CP violation is only observed in meson decays, and no conclusive signal of CP violation in baryon decays has been observed in particle physics experiments. The current observations of CP violation are not sufficient to explain the matter asymmetry of the universe, making the search for new sources of CP violation important.
Over the past 20 years, with the increasing number of measurements by BESIII, Belle, Belle II, and LHCb collaborations Belle:2004zjl ; Belle:2008xmh ; LHCb:2013xam ; LHCb:2015tgy ; BESIII:2015ysy ; BESIII:2015bjk ; LHCb:2017uwr ; BESIII:2017kqg ; Belle:2017ext ; LHCb:2018pcs ; LHCb:2018nfa , charmed baryons decays have attracted more attention from both the experimental and theoretical community Roberts:2007ni ; Briceno:2012wt ; Romanets:2012hm ; Brown:2014ena ; Lu:2016ogy ; He:2018php ; He:2018joe ; Zhao:2018mov ; He:2019tik ; Geng:2019xbo ; Jia:2019zxi , particularly on anti-triplet charmed baryon two-body decays. More precise experimental measurements will advance the search for CP violation hidden in baryon decays.
In the general theoretical analysis of CPV in Ref. PDG , the direct CPV of two-body decay processes is proportional to the weak phase difference and strong phase difference: $A_{CP}\propto\sin(\phi_{1}-\phi_{2})\sin(\delta_{1}-\delta_{2})$. Here, $\delta_{i}$ is the strong phase and $\phi_{i}$ is the weak phase from the CKM matrix element. For the charmed baryon two-body decays, the effective Hamiltonian $\mathcal{H}_{eff}$ is given as
$\mathcal{H}_{eff}$ $=$ $\frac{G_{F}}{\sqrt{2}}\bigg{(}\sum_{i=1,2}C_{i}\lambda O_{i}-\sum_{j=3}^{6}C_{j}\lambda_{b}O_{j}\bigg{)}+h.c.,$ (1)
where $O_{1,2}$ is the current-current operator and $O_{3\sim 6}$ is QCD penguin operator. $\lambda=V^{*}_{cq}V_{uq^{\prime}}$ with $q,q^{\prime}=d,s$ and $\lambda_{b}=V^{*}_{cb}V_{ub}$. The specific expressions of these operators refer to Ref. Buchalla:1995vs . Due to the different weak phases induced by the current-current operator and the penguin operator, the weak phase difference required by CPV is naturally generated. However, in previous studies He:2015fsa , the penguin operator is usually ignored owing to $|V^{*}_{cb}V_{ub}|\sim 10^{-4}<<|V^{*}_{cs}V_{ud}|$, so that CPV cannot be emerged.
To search for possible CPV in charmed baryon decays, we introduce SU(3) flavor symmetry in our work. The SU(3) flavor symmetry analysis, including the irreducible representation amplitude (IRA) and the topological diagrammatic approach (TDA), is one of the most powerful method in charmed baryon two-body decays. Previous studies show that SU(3) flavor symmetry in anti-triplet charmed baryon two-body decay is a good symmetry Xing:2023dni . In 2024, the BESIII Collaboration measured the decay branching ratios of $\Lambda^{+}_{c}$, and the Belle and Belle II Collaborations measured the decay branching ratios and asymmetry parameters of $\Xi_{c}^{0}$ for the first time BESIII:2023uvs ; Belle-II:2024jql ; BESIII:2024sfz as
$Br(\Lambda^{+}_{c}\to p\pi^{0})=(0.0156^{+0.0072}_{-0.0058}\pm 0.002)\%,$
$Br(\Lambda^{+}_{c}\to p\eta)=(0.163\pm 0.031\pm 0.011)\%,$
$Br(\Lambda^{+}_{c}\to pK_{L})=(1.67\pm 0.06\pm 0.04)\%,$
$Br(\Xi_{c}^{0}\to\Xi^{0}\pi^{0})=(0.69\pm 0.03\pm 0.05\pm 0.13)\%,$
$Br(\Xi_{c}^{0}\to\Xi^{0}\eta)=(0.16\pm 0.02\pm 0.02\pm 0.03)\%,$
$Br(\Xi_{c}^{0}\to\Xi^{0}\eta^{\prime})=(0.12\pm 0.03\pm 0.01\pm 0.02)\%,$
$\alpha(\Xi_{c}^{0}\to\Xi^{0}\pi^{0})=-0.90\pm 0.15\pm 0.23.$ (2)
Based on the new measurements, it is possible to determine all SU(3) IRA amplitudes and further understand the weak interactions of charmed particles, including possible CPV effects.
In our analysis, we provide a global fit to determine all the IRA amplitudes in Ref. Xing:2023dni . With the help of the equivalence between IRA and TDA, we can not only test the correctness of the topological diagrams but also estimate the value of each diagram in TDA. As a result, we find that although the topological diagrams in TDA are basically consistent with our IRA numerical results, the topological diagram which only includes the $O_{1,2}$ operator cannot perfectly explain our IRA results. Without loss of generality, we can use a new diagram to express the conflicting parts of IRA and TDA and include them in our fit without any assumptions about their source. To analyze the possible CPV introduced by the new diagram, we explored a novel way to determine their weak and strong phases. Our analysis shows that the new diagram may introduce a large CPV and it is promising to be observed in future experiments.
The global analysis of anti-triplet charmed baryon two-body decays with IRA- Using SU(3) flavor symmetry, the anti-triplet charmed baryon, light baryon octet, and light meson octet can be expressed by $3\times 3$ matrices $T_{c\bar{3}}$, $T_{8}$, and $P$, respectively Xing:2023dni . The SU(3) invariant decay amplitudes for anti-triplet charmed baryon two-body decays are constructed as $a_{15}\times\epsilon_{imn}(T_{c\bar{3}})^{[mn]}(H_{\overline{15}})^{\{ik\}}_{j}(\overline{T_{8}})^{j}_{k}P^{l}_{l}$ and $a_{6}\times(T_{c\bar{3}})^{[ik]}(H_{\overline{6}})_{\{ij\}}(\overline{T_{8}})^{j}_{k}P^{l}_{l}$. Here the effective weak interaction Hamiltonian with IRA method is decomposed: $3\otimes\bar{3}\otimes 3=3\oplus 3\oplus\bar{6}\oplus 15$. The nonzero matrix are $H_{\bar{6}}$ and $H_{15}$, which are expressed in Ref. Xing:2023dni . By enumerating all possible matrix combinations, nine independent amplitudes can be constructed as $\{a,b_{6},c_{6},d_{6}\}$ and $\{a^{\prime},b_{15},c_{15},d_{15},e_{15}\}$. Using these nine SU(3) amplitudes, all amplitudes of anti-triplet charmed baryon two-body decays can be expressed.
To correctly describe the phenomenological results of anti-triplet charmed baryon two-body decays, each SU(3) amplitude can be expressed by parity-violating form factors $f^{b,c,d}_{6,15},f^{a(\prime)}$ and parity-conserving form factors $g^{b,c,d}_{6,15},g^{a(\prime)}$ as defined in Ref. Xing:2023dni . Subsequently, the branching ratios and polarization parameters $\alpha$ are expressed in Ref. He:2015fsa .
By assuming real form factors for simplicity, we perform fits with updated data in Eq. 2 and the latest PDG PDG using the nonlinear least-$\chi^{2}$ method Fit . Here, we disregard the BESIII result $\alpha(\Lambda_{c}^{+}\to\Xi^{0}K^{+})=0.01\pm 0.16\pm 0.03$ BESIII:2023wrw because it has a large error compared to the central value and is inconsistent with our previous work Xing:2023dni . In addition the channel $\Xi_{c}^{0}\to\Xi^{0}\pi^{0}$ contributes the most to the $\chi^{2}$ while the channel $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ contributes less in the preliminary analysis. Therefore, we attempt to exclude this measurement from the global fit and obtain the fit results (Case I) in Table 1.
However, we predict $\alpha(\Lambda_{c}\to\Sigma^{0}K^{+})=-0.9971\pm 0.0036$ in this method, which deviates by $2\sigma$ from the experimental data $\alpha(\Lambda_{c}\to\Sigma^{0}K^{+})_{\text{exp}}=-0.54\pm 0.20$. This indicates that the assumption of real form factors is inadequate, prompting us to further explore the complex case (Case II) in the following.
The comprehensive analysis of anti-triplet charmed baryon two-body decays with the equivalence of TDA and IRA- All nine IRA amplitudes and 18 form factors can be determined based on the key experimental results measured by Belle and Belle II, as shown in Table 1. This facilitates an in-depth study of anti-triplet charmed baryon two-body decays. A comprehensive analysis can be provided by introducing the topological diagrammatic approach (TDA), which offers more intuitive physical images. In fact, the two different methods (IRA and TDA) are mutually equivalent, as studied in Refs. He:2018joe ; Hsiao:2021nsc ; Zhong:2024qqs . In our work, we use the same notation for TDA amplitudes ${\bar{a}_{1\sim 19}}$ in Ref. He:2018joe . The corresponding topological diagrams are shown in Fig. 1.
Using the definition $(T_{8})_{ijk}=\epsilon_{ijl}(\bar{T_{8}})^{l}_{k}$, the $H^{ij}_{k}$ in TDA can be decomposed into IRA forms as $H^{ij}_{k}=\frac{1}{2}\bigg{[}(H_{15})^{ij}_{k}+\frac{1}{2}\epsilon^{ijl}(H_{\bar{6}})_{kl}\bigg{]}$. The transformation from TDA to IRA occurs naturally with the following relations for the parameters:
$\overline{a}_{1}=b_{6}-d_{6}+e_{15},\quad\overline{a}_{2}=d_{6}-b_{6}+e_{15},\quad\overline{a}_{3}=-\frac{a}{2},$
$\overline{a}_{4}=\frac{1}{2}(-c_{6}+c_{15}),\quad\overline{a}_{5}=\frac{1}{2}a^{\prime},\quad\overline{a}_{7}=\frac{1}{2}(c_{6}+c_{15}),$
$\overline{a}_{6}=\frac{1}{2}(-b_{6}-c_{6}-e_{15}+d_{15})+\frac{1}{4}(a+a^{\prime}),$
$\overline{a}_{8}=\frac{1}{2}(b_{6}+c_{6}+d_{15}-e_{15})-\frac{1}{4}(a+a^{\prime}),$
$\overline{a}_{9}=\frac{1}{2}(a+a^{\prime}),\quad\overline{a}_{12}=c_{6},\quad\overline{a}_{15}=\frac{1}{2}(b_{6}-d_{6}+e_{15}),$
$\overline{a}_{10}=\frac{1}{2}(-b_{6}-c_{15}+d_{15}-e_{15})+\frac{1}{4}(a+a^{\prime}),$
$\overline{a}_{11}=\frac{1}{2}(b_{6}-c_{15}+d_{15}-e_{15})-\frac{1}{4}(a+a^{\prime})),$
$\overline{a}_{13}=\frac{1}{2}(-e_{15}+b_{15}+c_{6}+d_{15})-\frac{1}{4}(a^{\prime}-a)$
$\overline{a}_{14}=\frac{1}{2}(-e_{15}+b_{15}-c_{6}+d_{15})-\frac{1}{4}(a^{\prime}-a)$
$\overline{a}_{16}=\frac{1}{2}(-b_{6}+d_{6}+e_{15}),\quad\overline{a}_{17}=\frac{1}{2}(d_{6}+e_{15}-b_{15}),$
$\overline{a}_{18}=\frac{1}{2}(-d_{6}+e_{15}-b_{15}),\quad\overline{a}_{19}=d_{6}.$ (3)
By using the definition $\bar{a}_{i}=G_{F}\bar{u}(f_{i}-g_{i}\gamma_{5})u$ with $i={1\sim 19}$, the numerical results of the form factors for TDA are obtained in the lower panel of Table. 1.
Figure 1: Topology diagrams for the charmed baryon decays into an octet baryon and a light meson. Here $T_{1-7}$ show the Cabibbo-favored channels $c\to u\bar{d}s$. $T_{8,9}$ describe the new effects in the Cabibbo-suppressed $c\to u\bar{d}d(\bar{s}s)$ induced by $\textbf{f}_{b}$ and $\textbf{g}_{b}$ respectively.
Note that the diagrams $T_{1}-T_{2}$ and $T_{4}-T_{5}$ have the same topological structures, where one selects $\mathcal{M}_{T_{1}}=\mathcal{M}_{T_{2}}$ and $\mathcal{M}_{T_{4}}=\mathcal{M}_{T_{5}}$. The number of independent topological diagrams is reduced from seven to five, consistent with the latest TDA analysis in Ref. Zhong:2024qqs . As demonstrated in Ref. He:2018joe , each topological diagram in Fig. 1 corresponds to more than one TDA amplitude. Therefore, it is assumed that the absolute value of the TDA amplitude with the same topological diagram should be equal. By defining the form factor $\mathcal{M}_{T_{i}}=G_{F}\bar{u}(f_{T_{i}}-g_{T_{i}}\gamma_{5})u,\;i=1\sim 7$, one obtains the relations as
$|A_{1,15}|=A_{T_{1}}=|A_{2,16}|=A_{T_{2}},\quad|A_{3,5,9}|=A_{T_{3}},$
$|A_{4,6,10}|=A_{T_{4}}=|A_{7,8,11}|=A_{T_{5}},$
$|A_{12,13,14}|=A_{T_{6}},\quad|A_{17,18,19}|=A_{T_{7}},\;A=f,g.$ (4)
The assumptions are approximately supported by the numerical results within $1\sigma$ in Table. 1. However, some form factors contradict the assumption, such as $|f_{1,15}|-|f_{2,16}|\sim 0.06$, $|f_{T_{17,18}}|-|f_{19}|\sim 0.02$ and $|g_{17,19}|-|g_{18}|\sim 0.05$. To focus on the largest effects, we ignore the discrepancy involving $f_{19}$ and primarily consider $|f_{1,15}|>|f_{2,16}|$ and $|g_{17,19}|>|g_{18}|$ in the subsequent analysis. This implies that there are other effects in the topology diagrams. For exploring the possible new effects, one defines the form factors $\textbf{f}_{b}$ and $\textbf{g}_{b}$ to absorb these effect as $\textbf{f}_{b}=f_{T_{1}}-f_{T_{2}}$ and $\textbf{g}_{b}=-g_{18}-g_{T_{7}}$. The corresponding topology diagrams $T_{8,9}$ are given in Fig. 1. Then the IRA amplitude is expressed by TDA form factors and $\textbf{f}_{b},\textbf{g}_{b}$ as
$f^{a}=-\frac{f_{T_{6}}}{2},\quad f^{a\prime}=\frac{1}{2}f_{T_{6}}-f_{T_{3}},\quad f^{b}_{6}=\frac{3}{2}\textbf{f}_{b},$
$f^{b}_{15}=-f_{T_{6}}-f_{T_{7}},\quad f^{c}_{6}=f_{T_{6}}+f_{T_{4}},\quad f^{c}_{15}=f_{T_{4}},$
$f^{d}_{6}=-\frac{3}{2}\textbf{f}_{b}-f_{T_{7}},\quad f^{d}_{15}=-2f_{T_{4}}-f_{T_{6}},$
$f^{e}_{15}=-\frac{3}{2}\textbf{f}_{b}+3f_{T_{1}}+2f_{T_{4}}+f_{T_{6}}+f_{T_{7}},$
$g^{a}=g_{T_{4}}+\frac{g_{T_{6}}}{2},\quad g^{a\prime}=g_{T_{4}}+g_{T_{3}}-\frac{g_{T_{6}}}{2},\;g^{c}_{15}=0$
$g^{b}_{6}=-3g_{T_{1}}-2g_{T_{4}},\quad g^{b}_{15}=-g^{e}_{15}=\frac{\textbf{g}_{b}}{2}+g_{T_{6}}+g_{T_{7}},$
$g^{c}_{6}=g_{T_{6}},\quad g^{d}_{6}=3g_{T_{1}}-g_{T_{7}}+\frac{\textbf{g}_{b}}{2},\quad g^{d}_{15}=g_{T_{6}},$ (5)
Since all the diagrams induced by $O_{1,2}$ are included in TDA diagrams $T_{1-7}$, it is reasonable to believe the diagrams $T_{8,9}$ are induced by SU(3) symmetry breaking or other operators like the penguin operators $O_{3-6}$. The inclusion of these additional diagrams suggests the possibility of the involvement of a new weak phase in these processes. In Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, we find that $g^{c}_{15}=0$, which is consistent with our global fit (Case I) in Table. 1. One can also directly observe that $f^{b}_{6}$ depends only on new effect form factors, which provides us an ideal condition to study whether a different weak phase exists in these processes.
Table 1: The fit results for the real (Case I), complex (Case II) and TDA form factors.
form factors Case I ($\chi^{2}$/d.o.f=1.28)
vector(f) $f^{a}=0.0101(28)$ $f^{b}_{6}=0.0187(40)$ $f^{c}_{6}=0.0237(36)$ $f^{d}_{6}=-0.0093(37)$ $f^{a\prime}=-0.003(101)$
$f^{b}_{15}=-0.0100(34)$ $f^{c}_{15}=0.0073(34)$ $f^{d}_{15}=-0.0156(22)$ $f^{e}_{15}=0.0537(35)$
axial-vector(g) $g^{a}=-0.0266(78)$ $g^{b}_{6}=-0.1784(55)$ $g^{c}_{6}=0.0878(91)$ $g^{d}_{6}=-0.0556(72)$ $g^{a\prime}=0.05(44)$
$g^{b}_{15}=0.0746(48)$ $g^{c}_{15}=0.0075(90)$ $g^{d}_{15}=-0.0219(57)$ $g^{e}_{15}=0.0189(34)$
form factors Case II ($\chi^{2}$/d.o.f=0.597)
real part imaginary part
vector(f) $f^{a}=-0.0368(84)$ $f^{b}_{6}=0.0075(49)$ $f^{c}_{6}=0.017(13)$ $f_{T_{1}}=-0.0119(35)$ $f_{T_{4}}=0.0080(92)$
$f^{d}_{6}=0.006(27)$ $f^{a\prime}=-0.008(332)$ $f_{T_{6}}=-0.018(14)$ $f_{T_{7}}=0.014(13)$
$f^{b}_{15}=0.0201(71)$ $f^{c}_{15}=-0.0055(65)$ $f^{d}_{15}=-0.018(20)$ $\textbf{f}_{b}=-0.0013(58)$
$f^{e}_{15}=-0.0130(61)$
axial-vector(g) $g^{a}=-0.022(38)$ $g^{b}_{6}=0.035(15)$ $g^{c}_{6}=-0.036(21)$ $g_{T_{1}}=-0.024(51)$ $g_{T_{3}}=-0.18(60)$
$g^{d}_{6}=0.03(10)$ $g^{a\prime}=0.13(38)$ $g_{T_{4}}=0.11(12)$ $g_{T_{6}}=0.018(52)$
$g^{b}_{15}=-0.197(24)$ $g^{c}_{15}=-0.024(65)$ $g^{d}_{15}=-0.015(74)$ $g_{T_{7}}=0.02(10)$ weak phase
$g^{e}_{15}=-0.005(55)$ $\phi=-1.16(85)$
form factors TDA
vector(f) $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$
$f_{1}=0.0817(45)$ $f_{2}=0.0257(50)$ $f_{3}=-0.0051(14)$ $f_{4}=-0.0082(12)$ $f_{7}=0.0155(33)$
$f_{15}=0.0409(23)$ $f_{16}=0.0129(25)$ $f_{5}=-0.002(50)$ $f_{6}=-0.054(26)$ $f_{8}=-0.015(25)$
$f_{9}=0.003(50)$ $f_{10}=-0.046(26)$ $f_{11}=-0.031(24)$
$T_{6}$ $T_{7}$
$f_{12}=0.0237(36)$ $f_{17}=0.0272(18)$
$f_{13}=-0.024(25)$ $f_{18}=0.0365(35)$
$f_{14}=-0.048(24)$ $f_{19}=-0.0093(37)$
axial-vector(g) $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$
$g_{1}=-0.1039(86)$ $g_{2}=0.1416(77)$ $g_{3}=0.0133(39)$ $g_{4}=-0.0401(30)$ $g_{7}=0.0477(86)$
$g_{15}=-0.0519(43)$ $g_{16}=0.0708(38)$ $g_{5}=0.03(22)$ $g_{6}=0.03(11)$ $g_{8}=-0.07(11)$
$g_{9}=0.01(22)$ $g_{10}=0.07(11)$ $g_{11}=-0.12(11)$
$T_{6}$ $T_{7}$
$g_{12}=0.0878(91)$ $g_{17}=-0.0557(37)$
$g_{13}=0.04(11)$ $g_{18}=-0.00004(564)$
$g_{14}=-0.05(11)$ $g_{19}=-0.0556(72)$
The strong phase in charmed baryon two-body decays and CPV- Recently, numerous SU(3) analyses have introduced a strong phase in the form factors to resolve the puzzle of the BESIII measurement $\alpha(\Lambda_{c}^{+}\to\Xi^{0}K^{+})=0.01\pm 0.16\pm 0.03$ BESIII:2023wrw . In this section, we also analyze the situation using complex form factors (Case II) to conduct a comprehensive global analysis that incorporates all experimental data, including $\alpha(\Lambda_{c}^{+}\to\Xi^{0}K^{+})=0.01\pm 0.16\pm 0.03$ and $Br(\Xi_{c}^{0}\to\Xi_{0}\pi^{0})=(0.69\pm 0.03\pm 0.05\pm 0.13)\%$.
The above analysis demonstrates that the fits using the IRA method are largely consistent with the TDA assumption. Subsequently, we determine the strong phase from the perspective of topological diagrams. Since the strong phase arises from potential contributions of intermediate on-shell states in the decay process PDG , the imaginary part of the form factor in Eq. 4 corresponding to each topological diagram is incorporated into our global analysis. Thus, the imaginary part of the IRA form factor in our fitting is derived using Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis under the assumption: ${\mathcal{I}}m(\textbf{g}_{b})=0$. Additionally, we assume $\mathcal{I}m(f_{T_{3}})=0$ due to the global phase.
According to Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, the IRA form factor $f^{b}_{6}$ depends solely on the new effect, potentially introducing a new weak phase $\phi$ that could lead to significant CP violation.
In our global analysis, the form factors are defined as
$A^{q}_{6,15}$ $=$ $e^{i\phi_{1}}\bigg{(}\mathcal{R}e(A^{q}_{6,15})+\mathcal{I}m(A^{q}_{6,15})\bigg{)}=|A^{q}_{i}|e^{i\phi_{1}}e^{i\delta^{q}_{i}},$
$f^{b}_{6}$ $=$ $e^{i\phi}\bigg{(}\mathcal{R}e(f^{b}_{6})+\frac{3}{2}\mathcal{I}m(\textbf{f}_{b})\bigg{)}=|f^{b}_{6}|e^{i\phi}e^{i\delta^{b}_{6}},$
$f^{d}_{6}$ $=$ $e^{i\phi_{1}}\bigg{(}\mathcal{R}e(f^{d}_{6})+\mathcal{I}m(f^{d}_{6})+\mathcal{R}e(f^{b}_{6})\bigg{)}$
$-e^{i\phi}\bigg{(}\mathcal{R}e(f^{b}_{6})+\frac{3}{2}\mathcal{I}m(\textbf{f}_{b})\bigg{)},$
$f^{e}_{15}$ $=$ $e^{i\phi_{1}}\bigg{(}\mathcal{R}e(f^{e}_{15})+\mathcal{I}m(f^{e}_{15})+\mathcal{R}e(f^{b}_{6})\bigg{)}$ (6)
$-e^{i\phi}\bigg{(}\mathcal{R}e(f^{b}_{6})+\frac{3}{2}\mathcal{I}m(\textbf{f}_{b})\bigg{)},A=f,g,$
where $q=a,b,c,d,e$. The weak phase $\phi_{1}$ arises from the current-current operator as $\phi_{1}=\arg(V_{cq}^{*}V_{uq^{\prime}})\approx 0,-\pi$.
Although we can obtain the phase of $\textbf{f}_{b}$ in our fit, we cannot distinguish whether the phase is from the weak phase or the strong phase due to the lack of charmed anti-baryon decay data. Fortunately, in the SM, $\Delta c=1$ decay processes inducing CPV at the quark level are categorized into two types, each proportional to the CKM matrix elements $V_{ud}V_{cd}^{*}$ and $V_{us}V_{cs}^{*}$ He:2024pxh . This indicates that the weak phase only enters the Cabibbo-suppressed processes, which provides a new way to determine the strong phase and weak phase of $\textbf{f}_{b}$. Because SU(3) symmetry requires the strong phase in the same form factor corresponding to different processes to be equal Geng:2023pkr ; Zhong:2024qqs , one has the chance to determine the strong phase by the Cabibbo-allowed and doubly Cabibbo-suppressed processes. Based on the strong phase we derived, the weak phase can be further determined by the Cabibbo-suppressed processes.
By introducing the new weak phase $\phi$, one obtains the fit results with $\chi^{2}/d.o.f=0.597$, indicating that our fit is reasonable and the form factors effectively explain the experimental data.
With the determined weak phase $\phi=-1.16\pm 0.85$, one can predict the CP violation ($A_{CP}$) for the Cabibbo-suppressed processes. Given that the QCD penguin operator is likely responsible for this new weak phase, we observe that the fit vaule aligns with expectations for the QCD penguin’s weak phase. Hence, our study implies that the contribution of the QCD penguin may be more substantial than conventionally assumed, necessitating further theoretical investigation in the future.
Using the nontrivial weak phase $\phi$, we predict the values of $A_{CP}$ for the Cabibbo-suppressed channels and found that almost all the predicted $A_{CP}$ values are approximately zero within errors, except for the following processes
$A_{CP}^{\Lambda_{c}^{+}\to p\eta}=-0.047(45),\quad A_{CP}^{\Lambda_{c}^{+}\to n\pi^{+}}=-0.33(28),$
$A_{CP}^{\Xi_{c}^{+}\to\Xi^{0}K^{+}}=-0.39(32).$ (7)
We strongly recommend the experimental measurement of these two processes: $\Lambda_{c}^{+}\to p\eta/n\pi^{+}$ given their measured branching ratios. Additionally, the following processes also have a high probability of being observed in future experiments,
$A_{CP}^{\Xi_{c}^{+}\to\Sigma^{+}\eta}=-0.08(10),\quad A_{CP}^{\Xi_{c}^{+}\to\Lambda\pi^{+}}=-0.23(26),$
$A_{CP}^{\Xi_{c}^{0}\to\Sigma^{0}\pi^{0}}=-0.13(14).$ (8)
Conclusion- In our study, we investigate the equivalence of the SU(3) IRA and TDA methods through a global analysis (case I). Our findings indicate that the results obtained from the IRA method align well with the assumptions made in TDA, as illustrated by the topological diagram in Fig. 1. Nonetheless, certain form factors show discrepancies compared to the predicted values from the topological diagrams. To assess the extent of these discrepancies, we employ the form factors $\textbf{f}_{b}$ and $\textbf{g}_{b}$ to characterize the new effects.
By leveraging the equivalence between IRA and TDA methods, our global analysis extends to complex form factors. By determining the strong phase of new effect form factors $\textbf{f}_{b}$, the weak phase of this form factor can be determined in a completely new way with $\phi=-1.16\pm 0.85$. Then we further explore CP violation in anti-triplet charmed baryon two-body decay processes. Our analysis yields non-zero $A_{CP}$ in Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis and Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis. Considering its error, the observable CP violation can be observed at the order of $O(10^{-2})$. We strongly advocate for measuring these processes.
Note that the most ideal condition for measuring CPV is to produce charmed baryon and anti-baryon pairs simultaneously. Both the BESIII and Belle II Collaborations meet these conditions due to the most data on $\Lambda_{c}\bar{\Lambda}_{c}$ and $\Xi_{c}\bar{\Xi}_{c}$ pairs. For $\Lambda_{c}^{+}\to p\eta/n\pi^{+}$, their branching ratio has been measured by the BESIII and Belle II Collaborations. Therefore, it hold the potential for observing CP violation in charmed baryon decays at these two experimental facilities for the first time.
Acknowledgements- We thank Prof. Xiao-Gang He and Prof. Wei Wang for useful discussion. The work of Jin Sun is supported by IBS under the project code, IBS-R018-D1. The work of Zhi-Peng Xing is supported by NSFC under grant No.12375088 and No. 12335003. The work of Ruilin Zhu is supported by NSFC under grant No. 12322503 and No. 12075124, and by Natural Science Foundation of Jiangsu under Grant No. BK20211267.
Supplemental Material
Table 2: Experimental data and fitting results of anti-triplet charmed baryons two-body decays for two different fits. Case I(II) means the fit results for the real(complex) form factors.
channel exp Case I Case II
[…]
$\Lambda^{+}_{c}\to p\pi^{0}$ $0.0156(75)$ $0.0174(53)$ $0.0164(64)$
$\Lambda^{+}_{c}\to pK_{S}^{0}$ $1.59(7)$ $0.2(5)$ $1.646(48)$ $0.41(12)$ $1.634(49)$ $0.12(26)$
$\Lambda^{+}_{c}\to pK_{L}^{0}$ $1.67(7)$ $1.646(48)$ $1.634(49)$
$\Lambda^{+}_{c}\to p\eta$ $0.158(11)$ $0.158(11)$ $0.155(11)$
$\Lambda^{+}_{c}\to p\eta^{\prime}$ $0.048(9)$ $0.0464(61)$ $0.0541(76)$
$\Lambda^{+}_{c}\to\Lambda\pi^{+}$ $1.29(5)$ $-0.755(6)$ $1.305(47)$ $-0.7538(60)$ $1.290(49)$ $-0.7550(60)$
$\Lambda^{+}_{c}\to\Sigma^{0}\pi^{+}$ $1.27(6)$ $-0.466(18)$ $1.259(46)$ $-0.471(15)$ $1.261(49)$ $-0.472(15)$
$\Lambda^{+}_{c}\to\Sigma^{+}\pi^{0}$ $1.24(9)$ $-0.484(27)$ $1.273(46)$ $-0.469(15)$ $1.272(49)$ $-0.471(15)$
$\Lambda^{+}_{c}\to\Xi^{0}K^{+}$ $0.55(7)$ $0.01(16)$ $0.417(29)$ $0.549(69)$ 0.02(15)
$\Lambda^{+}_{c}\to\Lambda^{0}K^{+}$ $0.0642(31)$ $-0.58(5)$ $0.0641(29)$ $-0.556(44)$ $0.0643(31)$ $-0.579(50)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta$ $0.32(5)$ $-0.99(6)$ $0.327(49)$ $-0.9998(35)$ $0.321(50)$ $-0.973(57)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.41(8)$ $-0.46(7)$ $0.416(65)$ $-0.455(64)$ $0.354(66)$ $-0.446(66)$
$\Lambda^{+}_{c}\to\Sigma^{0}K^{+}$ $0.0370(31)$ $-0.54(20)$ $0.0376(18)$ $-0.9971(36)$ $0.0363(29)$ $-0.54(17)$
$\Lambda^{+}_{c}\to n\pi^{+}$ $0.066(13)$ $0.0641(23)$ $0.063(11)$
$\Lambda^{+}_{c}\to\Sigma^{+}K_{S}^{0}$ $0.047(14)$ $0.0311(29)$ $0.043(11)$
$\Xi^{+}_{c}\to\Xi^{0}\pi^{+}$ $1.6(8)$ $0.880(78)$ $1.90(68)$
$\Xi^{0}_{c}\to\Lambda K_{S}^{0}$ $0.32(6)$ $0.248(29)$ $0.331(59)$
$\Xi^{0}_{c}\to\Xi^{-}\pi^{+}$ $1.43(27)$ $-0.640(51)$ $1.17(18)$ $-0.701(45)$ $1.25(20)$ $-0.640(51)$
$\Xi^{0}_{c}\to\Xi^{-}K^{+}$ $0.039(11)$ $0.0520(80)$ $0.0485(74)$
$\Xi_{c}^{0}\to\Sigma^{0}K^{0}_{S}$ $0.054(16)$ $0.056(16)$ $0.055(16)$
$\Xi_{c}^{0}\to\Sigma^{+}K^{-}$ $0.18(4)$ $0.189(38)$ $0.182(40)$
$\Xi_{c}^{0}\to\Xi^{0}\pi^{0}$ $0.69(14)$ $-0.90(28)$ $0.153(48)$ $-0.45(13)$ $0.62(13)$ $-0.82(24)$
$\Xi_{c}^{0}\to\Xi^{0}\eta$ $0.16(4)$ $0.166(37)$ $0.16(4)$
$\Xi_{c}^{0}\to\Xi^{0}\eta^{\prime}$ $0.12(4)$ $0.125(38)$ $0.12(4)$
Table 3: The predicted values for branching ratios, polarization parameters ($\alpha,\beta,\gamma$) and CP violation with the final states $\eta^{(\prime)}$ for two different fits. Case I(II) means the fit results for the real(complex) form factors.
channel Case I Case II
[…]
$\Lambda^{+}_{c}\to\Sigma^{+}\eta$ $0.327(49)$ $-0.9998(35)$ $0.327(85)$ $-0.95(19)$ $-0.3(1.7)$ $0.1(3.3)$
$\Lambda^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.416(65)$ $-0.455(64)$ $0.38(10)$ $-0.442(82)$ $-0.70(97)$ $0.6(1.2)$
$\Lambda^{+}_{c}\to p\eta$ $0.154(10)$ $0.885(78)$ $0.162(23)$ $0.35(83)$ $-0.07(18)$ $-0.94(31)$ $-0.047(45)$
$\Lambda^{+}_{c}\to p\eta^{\prime}$ $0.0466(61)$ $-0.990(16)$ $0.059(20)$ $-0.82(63)$ $-0.6(1.1)$ $0.12(71)$ $-0.003(275)$
$\Xi^{+}_{c}\to\Sigma^{+}\eta$ $0.116(14)$ $0.91(11)$ $0.09(12)$ $0.3(1.0)$ $-0.1(1.1)$ $-0.93(23)$ $-0.08(10)$
$\Xi^{+}_{c}\to\Sigma^{+}\eta^{\prime}$ $0.117(16)$ $-0.418(71)$ $0.137(67)$ $-0.20(20)$ $-0.87(91)$ $0.4(1.8)$ $0.03(16)$
$\Xi^{+}_{c}\to p\eta$ $0.00908(67)$ $-0.054(61)$ $0.0107(70)$ $-0.15(21)$ $0.48(39)$ $-0.86(25)$
$\Xi^{+}_{c}\to p\eta^{\prime}$ $0.0091(10)$ $-0.9957(86)$ $0.0096(92)$ $-0.83(61)$ $-0.56(80)$ $-0.03(1.79)$
$\Xi^{0}_{c}\to\Xi^{0}\eta$ $0.177(75)$ $1.000(98)$ $0.082(68)$ $0.8(1.6)$ $0.5(9.8)$ $-0.3(14.9)$
$\Xi^{0}_{c}\to\Xi^{0}\eta^{\prime}$ $0.121(83)$ $1.0(2.2)$ $0.10(27)$ $-0.7(15.9)$ $-0.1(28.6)$ $-0.7(9.1)$
$\Xi^{0}_{c}\to\Sigma^{0}\eta$ $0.020(14)$ $0.7(2.0)$ $0.02(12)$ $-0.03(2.31)$ $-0.98(69)$ $-0.2(3.8)$ $-0.04(17)$
$\Xi^{0}_{c}\to\Sigma^{0}\eta^{\prime}$ $0.0070(74)$ $1.0(2.6)$ $0.010(24)$ $-0.7(8.6)$ $-0.2(13.2)$ $-0.6(5.3)$ $0.05(51)$
$\Xi^{0}_{c}\to\Lambda\eta$ $0.015(40)$ $0.9(2.6)$ $0.02(15)$ $-0.8(1.6)$ $-0.5(1.1)$ $-0.3(3.8)$ $-0.12(59)$
$\Xi^{0}_{c}\to\Lambda\eta^{\prime}$ $0.004(56)$ $0.4(23.2)$ $0.05(23)$ $-0.4(7.8)$ $0.8(5.4)$ $-0.5(16.5)$ $-0.2(1.01)$
$\Xi^{0}_{c}\to n\eta$ $0.0008(34)$ $-0.2(1.7)$ $0.0025(90)$ $-0.6(2.3)$ $-0.38(65)$ $-0.7(2.2)$
$\Xi^{0}_{c}\to n\eta^{\prime}$ $0.0005(52)$ $0.8(9.4)$ $0.001(15)$ $-0.2(4.3)$ $-0.04(19.52)$ $-0.99(13)$
Table 4: The predicted values for branching ratios, polarization parameters ($\alpha,\beta,\gamma$) and CP violation for two different fits. Case I(II) means the fit results for the real(complex) form factors.
channel Case I Case II
[…]
$\Lambda^{+}_{c}\to\Sigma^{0}\pi^{+}$ $1.259(46)$ $-0.471(15)$ $1.261(49)$ $-0.472(15)$ $0.78(11)$ $-0.42(20)$
$\Lambda^{+}_{c}\to\Lambda\pi^{+}$ $1.305(47)$ $-0.7538(60)$ $1.290(49)$ $-0.7550(60)$ $-0.47(12)$ $-0.46(12)$
$\Lambda^{+}_{c}\to\Sigma^{+}\pi^{0}$ $1.273(46)$ $-0.469(15)$ $1.272(49)$ $-0.471(15)$ $0.77(11)$ $-0.42(20)$
$\Lambda^{+}_{c}\to pK_{S}^{0}$ $1.646(48)$ $0.41(12)$ $1.634(49)$ $0.12(26)$ $0.26(39)$ $-0.96(13)$
$\Lambda^{+}_{c}\to\Xi^{0}K^{+}$ $0.417(29)$ $0.960(17)$ $0.549(69)$ $0.02(15)$ $-0.011(77)$ $0.9997(43)$
$\Xi^{+}_{c}\to\Sigma^{+}K_{S}^{0}$ $0.79(21)$ $0.51(18)$ $0.043(11)$ $-0.96(21)$ $0.3(1.1)$ $-0.08(1.28)$ $0.01(29)$
$\Xi^{+}_{c}\to\Xi^{0}\pi^{+}$ $0.880(78)$ $-0.892(39)$ $1.90(68)$ $-0.82(10)$ $0.47(18)$ $0.31(15)$
$\Xi^{0}_{c}\to\Sigma^{0}K_{S}^{0}$ $0.056(16)$ $-0.55(26)$ $0.055(16)$ $-0.2(3.7)$ $-0.5(3.3)$ $0.8(2.7)$
$\Xi^{0}_{c}\to\Lambda K^{0}_{S}$ $0.248(29)$ $0.87(15)$ $0.331(59)$ $-0.41(59)$ $-0.75(51)$ $-0.51(39)$
$\Xi^{0}_{c}\to\Sigma^{+}K^{-}$ $0.189(38)$ $0.91(25)$ $0.182(40)$ $0.4(2.0)$ $-0.18(62)$ $-0.91(93)$
$\Xi^{0}_{c}\to\Xi^{-}\pi^{+}$ $1.17(18)$ $-0.701(45)$ $1.25(20)$ $-0.640(51)$ $0.16(77)$ $-0.75(16)$
$\Xi^{0}_{c}\to\Xi^{0}\pi^{0}$ $0.153(48)$ $-0.45(13)$ $0.62(13)$ $-0.82(24)$ $-0.49(69)$ $-0.30(56)$
$\Lambda^{+}_{c}\to\Sigma^{0}K^{+}$ $0.0376(18)$ $-0.9971(36)$ $0.0363(29)$ $-0.54(17)$ $-0.07(39)$ $-0.84(11)$ $-0.02(10)$
$\Lambda^{+}_{c}\to\Lambda K^{+}$ $0.0641(29)$ $-0.556(44)$ $0.0643(31)$ $-0.579(50)$ $0.809(40)$ $-0.10(17)$ $-0.02(14)$
$\Lambda^{+}_{c}\to\Sigma^{+}K^{0}_{S}/K^{0}_{L}$ $0.0311(29)$ $-0.64(10)$ $0.043(11)$ $-0.96(21)$ $0.3(1.1)$ $-0.08(1.28)$ $0.01(29)$
$\Lambda^{+}_{c}\to p\pi^{0}$ $0.0174(53)$ $-0.10(13)$ $0.0164(64)$ $0.3(2.5)$ $-0.5(2.5)$ $0.79(59)$
$\Lambda^{+}_{c}\to n\pi^{+}$ $0.0641(23)$ $0.547(42)$ $0.063(11)$ $-0.56(21)$ $-0.81(15)$ $0.18(27)$ $-0.33(28)$
$\Xi^{+}_{c}\to\Sigma^{0}\pi^{+}$ $0.3175(88)$ $-0.731(17)$ $0.303(19)$ $-0.676(58)$ $0.56(11)$ $-0.48(12)$ $-0.010(23)$
$\Xi^{+}_{c}\to\Lambda\pi^{+}$ $0.0211(32)$ $-0.19(18)$ $0.047(18)$ $0.60(23)$ $-0.47(43)$ $-0.64(31)$ $-0.23(26)$
$\Xi^{+}_{c}\to\Sigma^{+}\pi^{0}$ $0.273(18)$ $0.43(12)$ $0.256(87)$ $-0.13(35)$ $0.43(54)$ $-0.89(31)$ $0.073(93)$
$\Xi^{+}_{c}\to pK^{0}_{S}/K^{0}_{L}$ $0.164(15)$ $-0.451(82)$ $0.178(90)$ $-0.86(28)$ $0.2(1.1)$ $-0.5(1.0)$ $0.007(169)$
$\Xi^{+}_{c}\to\Xi^{0}K^{+}$ $0.1336(54)$ $0.382(33)$ $0.107(23)$ $-0.48(20)$ $-0.70(16)$ $0.53(20)$ $-0.39(32)$
$\Xi^{0}_{c}\to\Sigma^{0}\pi^{0}$ $0.00002(12)$ $-0.99(74)$ $0.028(37)$ $-0.88(64)$ $-0.38(70)$ $-0.3(1.2)$ $-0.13(14)$
$\Xi^{0}_{c}\to\Lambda\pi^{0}$ $0.0331(47)$ $0.74(11)$ $0.0129(65)$ $-0.84(33)$ $0.25(68)$ $-0.47(44)$ $-0.03(26)$
$\Xi^{0}_{c}\to\Sigma^{+}\pi^{-}$ $0.0118(25)$ $0.89(27)$ $0.0118(26)$ $0.4(1.9)$ $-0.17(59)$ $-0.92(82)$
$\Xi^{0}_{c}\to pK^{-}$ $0.0146(39)$ $0.79(32)$ $0.0157(45)$ $0.3(1.6)$ $-0.14(50)$ $-0.95(56)$
$\Xi^{0}_{c}\to\Sigma^{-}\pi^{+}$ $0.0630(97)$ $-0.769(44)$ $0.082(13)$ $-0.577(97)$ $0.02(72)$ $-0.817(85)$ $-0.016(38)$
$\Xi^{0}_{c}\to nK^{0}_{S}/K^{0}_{L}$ $0.0234(54)$ $0.40(25)$ $0.022(45)$ $0.20(48)$ $0.01(46)$ $-0.979(99)$ $-0.14(33)$
$\Xi^{0}_{c}\to\Xi^{-}K^{+}$ $0.0520(80)$ $-0.656(44)$ $0.0485(74)$ $-0.69(10)$ $0.03(86)$ $-0.73(12)$ $-0.024(56)$
$\Xi^{0}_{c}\to\Xi^{0}K^{0}_{S}/K^{0}_{L}$ $0.0103(20)$ $0.59(33)$ $0.009(20)$ $0.31(70)$ $0.02(70)$ $-0.95(23)$ $-0.28(56)$
$\Lambda^{+}_{c}\to pK^{0}_{L}$ $1.743(51)$ $0.48(12)$ $1.73(13)$ $0.18(30)$ $0.23(40)$ $-0.96(14)$
$\Lambda^{+}_{c}\to nK^{+}$ $0.001009(90)$ $-0.975(21)$ $0.00227(79)$ $-0.86(10)$ $0.50(18)$ $0.06(17)$
$\Xi^{+}_{c}\to\Sigma^{0}K^{+}$ $0.01142(30)$ $-0.9964(13)$ $0.01048(74)$ $-0.661(60)$ $-0.06(16)$ $-0.748(49)$
$\Xi^{+}_{c}\to\Lambda K^{+}$ $0.00441(18)$ $0.615(30)$ $0.00432(87)$ $-0.50(16)$ $0.813(80)$ $0.31(27)$
$\Xi^{+}_{c}\to\Sigma^{+}K^{0}_{L}$ $0.98(23)$ $0.74(15)$ $0.8(1.3)$ $0.8(1.7)$ $-0.5(2.9)$ $0.2(1.2)$
$\Xi^{+}_{c}\to p\pi^{0}$ $0.00067(25)$ $0.06(32)$ $0.00070(72)$ $0.6(1.6)$ $0.88(89)$ $-0.2(2.4)$
$\Xi^{+}_{c}\to n\pi^{+}$ $0.00607(39)$ $0.939(21)$ $0.00410(52)$ $0.04(29)$ $-0.02(15)$ $0.999(15)$
$\Xi^{0}_{c}\to\Sigma^{0}K_{L}^{0}$ $0.070(18)$ $-0.27(28)$ $0.079(55)$ $-0.4(2.5)$ $-0.6(2.6)$ $0.7(3.3)$
$\Xi^{0}_{c}\to\Lambda K^{0}_{L}$ $0.230(26)$ $0.91(12)$ $0.316(72)$ $-0.44(65)$ $-0.76(48)$ $-0.47(32)$
$\Xi^{0}_{c}\to p\pi^{-}$ $0.00087(24)$ $0.77(33)$ $0.00095(29)$ $0.3(1.5)$ $-0.14(49)$ $-0.95(51)$
$\Xi^{0}_{c}\to\Sigma^{-}K^{+}$ $0.00285(44)$ $-0.736(45)$ $0.00339(53)$ $-0.607(50)$ $0.15(73)$ $-0.78(15)$
$\Xi^{0}_{c}\to n\pi^{0}$ $0.00162(25)$ $0.993(32)$ $0.00043(58)$ $-0.9(1.2)$ $0.14(47)$ $0.4(2.9)$